# Fine-tuning Neural Network Quantum States

### v2

@@ Abstract
We first discuss the results in one dimension. Here, the ground-state phase diagram shows two phases, separated by a Berezinskii-Kosterlitz-Thouless transition at $(J_{2}/J_{1})_{c}=0.24116(7)$ [32]: a gapless phase with no order whatsoever and a gapped one, with long-range dimer order. On finite clusters, the latter one may be extracted from the long-distance behavior of the dimer-dimer correlation functions ${D(r)=\braket{\hat{S}^{z}_{1}\hat{S}^{z}_{2}\hat{S}^{z}_{r}\hat{S}^{z}_{r+1}}-\braket{\hat{S}^{z}_{1}\hat{S}^{z}_{2}}\braket{\hat{S}^{z}_{r}\hat{S}^{z}_{r+1}}}$ [33, 34]. Specifically, performing a finite-size scaling, an estimation of the dimer order parameter can be obtained as $d^{2}=9|D(N/2-1)-2D(N/2)+D(N/2+1)|$ [33, 34]. However, we emphasize that the order parameter is exponentially small close to the transition, making it difficult to extract an accurate estimation of the actual value of $(J_{2}/J_{1})_{c}$ (indeed, the location of the transition may be easily obtained by looking at the level crossing between the lowest-energy triplet and singlet excitations [32]). As before, we pretrain at a given point, here $J_{2}/J_{1}=0.4$, and optimize the output layer of the network for different values of the frustrating ratio, both in the gapless and gapped regions. The results for $N=100$ (with periodic boundary conditions) are reported in Fig. 2b), again compared to DMRG calculations on the same cluster. In addition, in the inset of Fig. 2b), we compare the relative energy error $\Delta\varepsilon$ (with respect to the DMRG energies) of an RBM trained directly on the physical configurations [8] and of a fine-tuned ViT. This analysis underscores the importance of exploiting the features constructed by the pretrained ViT, resulting in an accuracy gain of more than two orders of magnitude with respect to the same network trained directly on configurations.
@@
Let us move on the discussion of how the output layer can modify the sign structure during the fine-tuning step. For the $J_{1}$-$J_{2}$ Heisenberg chain, the sign structure of the ground state wave function is not known except for $J_{2}=0$, where the so-called Marshall sign rule (MSR) [35] applies. However, even for large system sizes, the MSR constitutes an accurate approximation of the sign structure up to $J_{2}/J_{1}\leq 0.5$ [36]. In Fig. 4, we show the predicted phases (00 or $\pi$), on top of the UMAP projections of the vectors $\bm{z}_{p}$ generated by the pretrained network at $J_{2}/J_{1}=0.4$. At $J_{2}/J_{1}=0.1$ (see the left panel of Fig. 4), after the fine-tuning procedure, the signs exactly match the ones obtained at $J_{2}/J_{1}=0.4$ (not shown). This is because, at the pretraining point, where the clusters are formed, the MSR remains a highly accurate approximation of the ground state sign structure. By contrast, for $J_{2}/J_{1}=0.6$, this is no longer true, and the output layer must adjust the phases accordingly (see the right panel of Fig. 4); still, the fine-tuned ViT performs better than a RBM trained on spin configurations, refer to Supplemental Material .3 for a detailed discussion.
@@
Finally, we consider the two-dimensional $J_{1}$-$J_{2}$ Heisenberg model on an $L\times L$ square lattice. The ground state of this model features magnetic order in the two limits $J_{1}\ll J_{2}$ and $J_{1}\gg J_{2}$. Its presence can be characterized with the spin structure factor ${S(\bm{k})=\sum_{\bm{R}}e^{i\bm{k}\cdot\bm{R}}\braket{\hat{\bm{S}}_{\bm{0}}\cdot\hat{\bm{S}}_{\bm{R}}}}$, where $\bm{R}$ runs over all the lattice sites of the square lattice. Specifically, when $J_{2}=0$ the model reduces to the unfrustrated Heisenberg model where long-range Néel order is present [37, 38]. The latter one can be detected by measuring ${m^{2}_{\text{Néel}}=S(\pi,\pi)/L^{2}}$. In the opposite regime $J_{2}/J_{1}\rightarrow\infty$, the system exhibits instead columnar magnetic order, identified by the order parameter ${m^{2}_{\text{stripe}}=[S(0,\pi)+S(\pi,0)]/(2L^{2})}$. In the intermediate region, around $J_{2}/J_{1}\approx 0.5$ the system is highly frustrated and the nature of the ground state is still under debate [10, 39, 40, 41, 42]. Here, we limit ourselves to the $6\times 6$ cluster, where exact diagonalizations are possible (no DMRG calculations on the structure factor are available on larger clusters with periodic boundary conditions). We first perform the pretraining at $J_{1}/J_{2}=0.5$, then perform the fine-tuning for $10.2<J_{2}/J_{1}<$ 1 and evaluate the order parameters $m^{2}_{\text{Néel}}$ and $m^{2}_{\text{stripe}}$, see in Fig. 2c). Remarkably, even for this complicated two-dimensional model, the correct behavior of the two magnetic order parameters can be reconstructed with great accuracy starting from a single pretrained ViT.
@@
Conclusions. We showed that, for several physical systems exhibiting phase transitions, pretraining a Neural Network Quantum State near the transition point yields a set of features that can be fine-tuned to obtain accurate descriptions of the phase diagram. In addition, the analysis of the feature space facilitated the extraction of valuable insights into the structure of the wave function, enhancing the interpretability of NQS. Furthermore, the fine-tuning process proves to be computationally more efficient, both in terms of time and memory, compared to traditional approaches in which the full training is required at each point of the phase diagram. We contend that, akin to the prevalent approach in machine learning [43], the employment of pretrained networks holds great potential for advancing the exploration of physical systems through NQS. This aligns with the findings of a recent study by Ref. [44], demonstrating that, in electronic-structure problems, a single wave function can be employed to investigate multiple compounds and geometries. A similar framework has also been proposed for decoding quantum error-correcting codes with generative modeling [45]. Exploring extensions of this approach to perform fine-tuning across different physical models stands as a crucial topic for future studies, as well as the development of techniques to automatically identify the most expressive pretraining point [46, 47, 48, 49]. Another possible application could focus on approximating the real-time dynamics [50, 51] by adjusting only the parameters of the shallow output network, thereby solving the problem within the feature space rather than the configuration space.
@@ Acknowledgements.
SUPPLEMENTAL MATERIAL
.1 Memory Efficiency in Fine-tuning and Pretraining Processes
The primary constraint in training neural networks with a large number of parameters arises from the restricted memory capacity of contemporary graphical processing units (GPUs), rather than their computational speed. Specifically, this limitation is associated to the back-propagation algorithm [1], that is crucial for evaluating the gradients of the network efficiently, but whose memory cost scales with the depth of the computation.
Consider a deep neural network that takes an input vector $\bm{x}$ and produces a scalar output $f(\bm{x},\theta)\in\mathbb{R}$, where $\theta$ is a vector of trainable parameters. For simplicity, we arrange these parameters as $\theta=\text{Concat}(\theta_{0},\dots,\theta_{n_{l}})$, where $\theta_{l}$ is a vector containing all the $P_{l}$ parameters of the $l$-th layer, and $P$ is the total number of parameters across all layers, i.e., $P=\sum_{l=1}^{n_{l}}P_{l}$. Additionally, assume that, when computing the output, the network generates $K$ intermediate activations $a_{k}$, each of size $A_{k}$, and $A$ is the overall number of activations calculated as $A=\sum_{k=1}^{K}A_{k}$. For a batch of $M$ distinct input vectors, the loss function can be defined as $\mathcal{L}(\theta)=(1/M)\sum_{i=1}^{M}L[f(\bm{x}_{i},\theta)]$. To efficiently backpropagate the gradients of the loss with respect to the parameters, it is necessary to store all the $A$ activations. Thus the total memory cost of the algorithm scales with the depth of the computations and is expressed as $M\times(A+\text{max}_{l}P_{l})$ (neglecting the cost of storing all $P$ weights). On the contrary, for the forward pass the memory cost is independent of the computation depth and is equal to $M\times(\text{max}_{k}A_{k}+\text{max}_{l}P_{l})$. Further details can be found in Ref. [53].
Notice that, during the fine-tuning process, the memory-intensive backward pass over the deep network becomes unnecessary. In the context of this paper, for the used ViT architectures, the memory needed during the fine-tuning stage is approximately ten times less than what is required during the pretraining stage.
Figure 5: Relative error $\Delta\varepsilon$ of the energy with respect to DMRG for the $J_{1}$-$J_{2}$ Heisenberg model on a chain [see Eq. (4)] of $N=100$ sites. The curves are obtained performing the fine-tuning procedure starting from different pretraining points generated by a ViT with hyperparameters $h=12$, $d=192$, $n_{l}=4$. Specifically, we set $J_{2}/J_{1}=0.4$ (green triangles), $J_{2}/J_{1}=0.1$ (blue circles), $J_{2}/J_{1}=0.6$ (red squares). The accuracy of the same fully-connected network (RBM) optimized on the physical configurations is also reported for comparison (orange diamonds).
.2 Choosing different pretraining points
@@
The accuracy of the fine-tuning across various points on the phase diagram is influenced by the choice of the pretraining point. In our study, we have always pretrained near transition points, where we expect better generalization properties as discussed in the main text. Here, we investigate how the accuracy of the fine-tuned results varies when choosing different pretraining points, for example within the bulk of one phase. In Fig. 5, we show the accuracy of the energy $\Delta\varepsilon$ relative to DMRG calculations for the $J_{1}$-$J_{2}$ Heisenberg model on a chain [see Eq. (4)] of $N=100$ sites (assuming periodic boundary conditions). The transition point of the model in the thermodynamic limit is $(J_{2}/J_{1})_{c}=0.24116(7)$; however, on a finite system with $N=100$ sites, the point exhibiting the maximum slope in the dimer order parameter occurs around $J_{2}/J_{1}=0.4$ (refer to the central panel of Fig. 2). The accuracy of the fine-tuned energies, using $J_{2}/J_{1}=0.4$ as the pretraining point, is approximately $\Delta\varepsilon\approx 10^{-3}$ within the interval $J_{2}/J_{1}\in[0.1,0.6]$ (green triangles in Fig. 5). Conversely, pretraining from $J_{2}/J_{1}=0.1$ (blue circles in Fig. 5) yields higher accuracy before $J_{2}/J_{1}=0.4$, but as the distance from the pretraining point increases the accuracy deteriorates to approximately $\Delta\varepsilon\approx 10^{-2}$. A similar behaviour can be observed when choosing $J_{2}/J_{1}=0.6$ as the pretraining point (red squares in Fig. 5). Consequently, selecting a pretraining point that lies near the transition appears to strike the optimal balance, yielding an accuracy roughly consistent across all other points within the phase diagram. Nevertheless, it is worth noting that training the RBM in the hidden space consistently outperforms direct training on physical configurations, as illustrated by the orange diamonds in Fig. 5.
.3 Comparison with Marshall Sign Prior
A relevant question is trying to understand which kind of prior information is encoded in the features generated by the pretrained network.

### v3

@@ III.2 The one-dimensional $J_{1}$-$J_{2}$ Heisenberg model
We first discuss the results in one dimension. Here, the ground-state phase diagram shows two phases, separated by a Berezinskii-Kosterlitz-Thouless transition at $(J_{2}/J_{1})_{c}=0.24116(7)$ [34]: a gapless phase with no order whatsoever and a gapped one, with long-range dimer order. On finite systems, the latter one may be extracted from the long-distance behavior of the dimer-dimer correlation functions ${D(r)=\braket{\hat{S}^{z}_{1}\hat{S}^{z}_{2}\hat{S}^{z}_{r}\hat{S}^{z}_{r+1}}-\braket{\hat{S}^{z}_{1}\hat{S}^{z}_{2}}\braket{\hat{S}^{z}_{r}\hat{S}^{z}_{r+1}}}$ [35, 36]. Specifically, performing a finite-size scaling, an estimation of the dimer order parameter can be obtained as $D^{2}=9|D(N/2-1)-2D(N/2)+D(N/2+1)|$ [35, 36]. However, we emphasize that the order parameter is exponentially small close to the transition, making it difficult to extract an accurate estimation of the actual value of $(J_{2}/J_{1})_{c}$ (indeed, the location of the transition may be easily obtained by looking at the level crossing between the lowest-energy triplet and singlet excitations [34]). As before, we pretrain at a given point, here $J_{2}/J_{1}=0.4$, and optimize the output layer of the network for different values of the frustrating ratio, in both the gapless and gapped regions. The results for $N=100$ (with periodic boundary conditions) are reported in Fig. 2b), again compared to DMRG calculations on the same system. In addition, in the inset of Fig. 2b), we compare the relative energy error $\Delta\varepsilon$ (with respect to the DMRG energies) of an RBM trained directly on the physical configurations [8] and of a fine-tuned ViT. This analysis underscores the importance of exploiting the features constructed by the pretrained ViT, resulting in an accuracy gain of more than two orders of magnitude with respect to the same network trained directly on configurations.
$\bm{J_{2}/J_{1}}$ DMRG ViT Fine-tuning
0.10 -0.425417395 -0.4254174 -0.425218
0.20 -0.408572967 -0.4085728 -0.408453
0.30 -0.393126745 -0.3931204 -0.393059
0.40 -0.380387370 -0.3803726 -0.380370
0.60 -0.380804138 -0.3807913 -0.379902
Table 1: Variational ground-state energies for the $J_{1}$-$J_{2}$ Heisenberg chain with system size $N=100$. The DMRG computations are conducted employing a bond dimension up to $\chi=10^{3}$ under periodic boundary conditions. For both instances involving ViT, one trained from scratch and another pretrained at $J_{2}/J_{1}=0.4$ followed by fine-tuning, the Monte Carlo error attributed to finite sampling effects affects the last digit of the reported results.
For completeness, we report the ground-state energies for various frustration ratios $J_{2}/J_{1}$ in Table 1. They are obtained through three distinct methodologies: DMRG, ViT trained from scratch, and ViT pretrained at ${J_{2}/J_{1}=0.4}$ and subsequently fine-tuned for other frustration ratios. Notably, the fine-tuned ViT exhibits remarkable accuracy when compared to DMRG results, reaching an accuracy $\Delta\varepsilon\lesssim 10^{-3}$ for all the values of the frustration ratio in the interval $J_{2}/J_{1}\in[0.1,0.6]$.
Figure 4: Graphical representation of the hidden representations for the $J_{1}$-$J_{2}$ Heisenberg chain. The data points, corresponding to a sample of $M=3000$ physical configurations, represent UMAP projections of vectors $\bm{z}_{p}$ generated by a ViT with hyperparameters $h=12$, $d=192$, and $n_{l}=4$, pretrained at the point $J_{2}/J_{1}=0.4$ for a system size of $N=100$. The depicted colors correspond to the predicted phases (00 or $\pi$) after fine-tuning at two specific points within the phase diagram: $J_{2}/J_{1}=0.1$ [panel (a)] and $J_{2}/J_{1}=0.6$ [panel (b)]. Panel (a) reveals a close resemblance between the cluster structure identified during the pretraining at $J_{2}/J_{1}=0.4$, which matches the Marshall Sign Rule.
@@
Let us move to the discussion of how the output layer can modify the sign structure during the fine-tuning step. For the $J_{1}$-$J_{2}$ Heisenberg chain, the sign structure of the ground-state wave function is not known except for $J_{2}=0$, where the so-called Marshall Sign Rule (MSR) [37] applies. However, even for large system sizes, the MSR constitutes an accurate approximation of the sign structure up to $J_{2}/J_{1}\leq 0.5$ [38]. In Fig. 4, we show the predicted phases (00 or $\pi$), on top of the UMAP projections of the vectors $\bm{z}_{p}$ generated by the pretrained network at $J_{2}/J_{1}=0.4$. At $J_{2}/J_{1}=0.1$ [see Fig. 4(a)], after the fine-tuning procedure, the signs exactly match the ones obtained at $J_{2}/J_{1}=0.4$ (not shown). This is because, at the pretraining point, where the clusters are formed, the MSR remains a highly accurate approximation of the ground-state sign structure. By contrast, for $J_{2}/J_{1}=0.6$, this is no longer true, and the output layer must adjust the phases accordingly [see Fig. 4(b)]; still, the fine-tuned ViT performs better than a RBM trained on spin configurations; see Appendix D for a detailed discussion.
III.3 The two-dimensional $J_{1}$-$J_{2}$ Heisenberg model
@@
Finally, we consider the two-dimensional $J_{1}$-$J_{2}$ Heisenberg model on an $L\times L$ square lattice. The ground state of this model features magnetic order in the two limits $J_{1}\ll J_{2}$ and $J_{1}\gg J_{2}$. Its presence can be characterized with the spin structure factor ${S(\bm{k})=\sum_{\bm{R}}e^{i\bm{k}\cdot\bm{R}}\braket{\hat{\bm{S}}_{\bm{0}}\cdot\hat{\bm{S}}_{\bm{R}}}}$, where $\bm{R}$ runs over all the lattice sites of the square lattice. Specifically, when $J_{2}=0$ the model reduces to the unfrustrated Heisenberg model where long-range Néel order is present [39, 40]. The latter one can be detected by measuring ${m^{2}_{\text{Néel}}=S(\pi,\pi)/L^{2}}$. In the opposite regime $J_{2}/J_{1}\rightarrow\infty$, the system exhibits instead columnar magnetic order, identified by the order parameter ${m^{2}_{\text{stripe}}=[S(0,\pi)+S(\pi,0)]/(2L^{2})}$. In the intermediate region, around $J_{2}/J_{1}\approx 0.5$ the system is highly frustrated and the nature of the ground state is still under debate [10, 41, 42, 43, 44]. Here, we limit ourselves to the $6\times 6$ system, where exact diagonalizations are possible (no DMRG calculations on the structure factor are available on larger systems with periodic boundary conditions). We first perform the pretraining at $J_{1}/J_{2}=0.5$, then perform the fine-tuning for $10.2<J_{2}/J_{1}<$ 1 and evaluate the order parameters $m^{2}_{\text{Néel}}$ and $m^{2}_{\text{stripe}}$, see Fig. 2c). Remarkably, even for this complicated two-dimensional model, the correct behavior of the two magnetic order parameters can be reconstructed with great accuracy starting from a single pretrained ViT.
IV Conclusions
@@
We showed that, for several physical systems exhibiting phase transitions, pretraining a Neural-Network Quantum State near the transition point yields a set of features that can be fine-tuned to obtain accurate descriptions of the phase diagram. In addition, the analysis of the feature space facilitated the extraction of valuable insights into the structure of the wave function. Furthermore, the fine-tuning process proves to be computationally more efficient, in terms of both time and memory, compared to traditional approaches in which the full training is required at each point of the phase diagram. We contend that, akin to the prevalent approach in machine learning [45], the employment of pretrained networks holds great potential for advancing the exploration of physical systems through NQS. This aligns with the findings of a recent study by Ref. [46], demonstrating that, in electronic-structure problems, a single wave function can be employed to investigate multiple compounds and geometries. A similar framework has also been proposed for decoding quantum error-correcting codes with generative modeling [47]. Exploring extensions of this approach to perform fine-tuning across different physical models stands as a crucial topic for future studies, as well as the development of techniques to automatically identify the most expressive pretraining point [48, 49, 50, 51]. Another possible application could focus on approximating the real-time dynamics [52, 53] by adjusting only the parameters of the shallow output network, thereby solving the problem within the feature space rather than the configuration space. It would be intriguing to further explore the precise physical meaning of the learned features, such as establishing a connection between the clusters in the feature space and the order parameters.
@@ Acknowledgements.
Appendix A Choosing different pretraining points
@@
The accuracy of the fine-tuning across various points on the phase diagram is influenced by the choice of the pretraining point. In our study, we have always pretrained near transition points, where we expect better generalization properties as discussed in the main text. Here, we investigate how the accuracy of the fine-tuned results varies when choosing different pretraining points, for example within the bulk of one phase. In Fig. 5, we show the accuracy of the energy $\Delta\varepsilon$ relative to DMRG calculations for the $J_{1}$-$J_{2}$ Heisenberg model on a chain [see Eq. (4)] of $N=100$ sites (assuming periodic boundary conditions). The transition point of the model in the thermodynamic limit is $(J_{2}/J_{1})_{c}=0.24116(7)$; however, on a finite system with $N=100$ sites, the point exhibiting the maximum slope in the dimer order parameter occurs around $J_{2}/J_{1}=0.4$ (refer to the central panel of Fig. 2). The accuracy of the fine-tuned energies, using $J_{2}/J_{1}=0.4$ as the pretraining point, is approximately $\Delta\varepsilon\approx 10^{-3}$ within the interval $J_{2}/J_{1}\in[0.1,0.6]$ (green triangles in Fig. 5). Conversely, pretraining from $J_{2}/J_{1}=0.1$ (blue circles in Fig. 5) yields higher accuracy before $J_{2}/J_{1}=0.4$, but as the distance from the pretraining point increases the accuracy deteriorates to approximately $\Delta\varepsilon\approx 10^{-2}$. A similar behavior can be observed when choosing $J_{2}/J_{1}=0.6$ as the pretraining point (red squares in Fig. 5).
It is interesting to note that the accuracy of the network pretrained at $J_{2}/J_{1}=0.1$ (blue circles) deteriorates by four orders of magnitude when fine-tuning at $J_{2}/J_{1}=0.4$. In contrast, the network pretrained at $J_{2}/J_{1}=0.4$ (green triangles) loses less than one order of magnitude in accuracy when finetuning at $J_{2}/J_{1}=0.1$, with the error rising from $\Delta\epsilon\approx 10^{-4}$ to $\Delta\epsilon\approx 10^{-3}$. This result suggests that features learned near the phase transition are more robust for generalization compared to those learned within the bulk of a phase. Consequently, selecting a pretraining point that lies near the transition appears to strike the optimal balance, yielding an accuracy roughly consistent across all other points within the phase diagram. Nevertheless, it is worth noting that training the RBM in the hidden space consistently outperforms direct training on physical configurations, as illustrated by the orange diamonds in Fig. 5.
Figure 5: Relative error $\Delta\varepsilon$ of the energy with respect to DMRG for the $J_{1}$-$J_{2}$ Heisenberg model on a chain [see Eq. (4)] of $N=100$ sites. The curves are obtained performing the fine-tuning procedure starting from different pretraining points generated by a ViT with hyperparameters $h=12$, $d=192$, $n_{l}=4$. Specifically, we set $J_{2}/J_{1}=0.4$ (green triangles), $J_{2}/J_{1}=0.1$ (blue circles), and $J_{2}/J_{1}=0.6$ (red squares). The accuracy of the same fully connected network (RBM) optimized on the physical configurations is also reported for comparison (orange diamonds).
Appendix B Memory Efficiency in Fine-tuning and Pretraining Processes
The primary constraint in training neural networks with a large number of parameters arises from the restricted memory capacity of contemporary graphical processing units (GPUs), rather than their computational speed. Specifically, this limitation is associated to the back-propagation algorithm [1], which crucial for evaluating the gradients of the network efficiently, but whose memory cost scales with the depth of the computation. Consider a deep neural network that takes an input vector $\bm{x}$ and produces a scalar output $f(\bm{x},\theta)\in\mathbb{R}$, where $\theta$ is a vector of trainable parameters. For simplicity, we arrange these parameters as $\theta=\text{Concat}(\theta_{0},\dots,\theta_{n_{l}})$, where $\theta_{l}$ is a vector containing all the $P_{l}$ parameters of the $l$-th layer, and $P$ is the total number of parameters across all layers, i.e., $P=\sum_{l=1}^{n_{l}}P_{l}$. Additionally, assume that, when computing the output, the network generates $K$ intermediate activations $a_{k}$, each of size $A_{k}$, and $A$ is the overall number of activations calculated as $A=\sum_{k=1}^{K}A_{k}$. For a batch of $M$ distinct input vectors, the loss function can be defined as $\mathcal{L}(\theta)=(1/M)\sum_{i=1}^{M}L[f(\bm{x}_{i},\theta)]$. To efficiently backpropagate the gradients of the loss with respect to the parameters, it is necessary to store all the $A$ activations. Thus the total memory cost of the algorithm scales with the depth of the computations and is expressed as $M\times(A+\text{max}_{l}P_{l})$ (neglecting the cost of storing all $P$ weights). On the other hand, for the forward pass the memory cost is independent of the computation depth and is equal to $M\times(\text{max}_{k}A_{k}+\text{max}_{l}P_{l})$. Further details can be found in Ref. [55]. Notice that, during the fine-tuning process, the memory-intensive backward pass over the deep network becomes unnecessary. In the context of this paper, for the used ViT architectures, the memory needed during the fine-tuning stage is approximately ten times less than what is required during the pretraining stage. The backpropagation of gradients constitutes the primary memory bottleneck, even when employing the Stochastic Reconfiguration optimization method [21]. This method requires the allocation of a matrix containing $4M^{2}$ real numbers, where $M$ denotes the number of samples used in optimization. With double precision, this memory requirement translates to $32M^{2}/10^{9}$ GB. In our optimizations, with $M=3000$, the memory usage is approximately 0.3 GB. This is two orders of magnitude smaller than the memory required for the backward pass during pretraining.
Figure 6: Correlation between the local magnetization $\sum_{i=1}^{N}\sigma_{i}$ and the principal component of the hidden representations of the configurations associated to the ViT used to obtain the results for the Ising model in transverse field (see Results section). In the inset, the PCA spectrum is shown.